### v3

@@ 1 Introduction
As a crucial information carrier, underwater images play an important role in ocean tasks[1]. Unfortunately, underwater images often encounter various degradation problems due to two main factors [2]. Firstly, different wavelengths of light decay at varying rates in the underwater medium, resulting in a prevalent blue-green color deviation. Secondly, the scattering effect of suspended particles hinders normal light propagation, leading to problems such as low contrast and blurred details [3]. Meanwhile, these degradations worsen with the increasing distance from the imaging scene to the camera [4]. Therefore, underwater image enhancement (UIE) methods have been widely applied to improve vision quality, highlighting important practical value.
Figure 1: Near-far scenarios enhancement comparison by different methods, where the orange box represents the enlarged near scene and the red box represents the enlarged far scene. The top row shows the raw image, the results of L2UWE [5], CBM [6], TEBCF [7], WWPE [8], and Water-Net [9]. The bottom row shows the results of SCNet [10], TACL [11], USUIR [12], PUGAN [13], Spectroformer [14], and our UVZ.
In the early days of UIE field, researchers have proposed model-free and model-based traditional methods. Model-free methods [5, 15, 16] enhances underwater images by modifying pixels in degraded channels and blurred regions. On the other hand, model-based methods [7, 17] introduce physical cues to build the underwater degradation process, and reverse it to restore underwater images. In recent years, deep learning methods [11, 18, 19] have shown remarkable enhancements, leveraging benchmark data and miscellaneous networks. Despite significant progress achieved with the above methods, challenges emerge when addressing widely varying degrees of regional degradation. As shown in Fig. 1, most UIE methods have considerable visual improvement in the near scene, but still retain notable color deviation in the far scene. Fig. 2 highlights the significant differences between near and far scenes, thereby underscoring the motivation behind this work. A more detailed description is provided in Section 3.1.
@@
For more precise color enhancement, we believe that the integration of depth information related to scene distance is essential, which benefits the region division in underwater images. Previous works [12, 13, 20] attempted to incorporate depth information related to scene distance, contributing to the response of degraded regions. However, direct weight adjustments may restrict the exploration of visual cues, as the adjacent features of far and near points exhibit varying gains in the enhancement process. In near regions of relatively good quality, local information from surrounding features is critical to the quality of the enhanced image, whereas in the far regions of heavily degradation, non-local information from other regions can provide more valuable insights. With the above analysis, we aim to construct an adaptive region enhancement method in underwater scenes, achieving comprehensive visual enhancement by differentiating in utilizing local and non-local information. To achieve this goal, there are two main challenges:(1) On the one hand, the local feature capture capability helps to enhance the near region where the surrounding information is enough. On the other hand, the remote dependency establishment capability helps to enhance the far region where the surrounding information is lost. However, due to feature space differences, there is mutual interference between these two capabilities. (2) Diverse underwater scenes tremendously raise the difficulty of predicting the depth map.
@@
• We design a novel depth-guided perception framework, UVZ. For non-homogeneous degraded underwater scenes, UVZ employs the adapted depth maps to implement a learning strategy that adaptively incorporates non-local and local features.
• To enhance the ability to distinguish degraded regions, we incorporate an attention estimation-scene reconstruction training process, aiming to strengthen the connection between depth maps and underwater scenes.
• Compared to state-of-the-art methods, UVZ demonstrates more attractive results across extensive qualitative and quantitative comparisons. Moreover, UVZ exhibits generalization abilities to other visual tasks without parameter tuning.
@@ 2.1 Traditional Methods
In response to the degradation characteristics of images, model-free methods notably enhance the contrast and saturation of the images. The early studies focused on adjusting image pixel values in the spatial domain. Ancuti et al. [15] fused color compensated and white balance images to obtain natural underwater image. Marques et al. [5] used contrast to guide the generation of images with salient details and brightness enhancement. As the research progressed, some methods introduced morphology and structural properties to understand underwater images. Based on the concept of contour bougie morphology, Yuan et al. [6] enhanced the image quality by mitigating scattering blur and contrast stretching. Kang et al. [21] decompose the image into mean intensity, contrast, and structure, the three components are fused in different ways to obtain the enhancement result. Additionally, Zhang et al. [3] introduced the attenuation characteristics to correct the color, and introduced the integral and squared integral maps to enhance the contrast. Overall, model-free methods overlook the characteristics of underwater degradation process, especially the depth between the object and the camera [22, 23], leading to color distortion or improper enhancement.
Model-based methods explore the process of underwater image degradation and use priors and assumptions to model this process. Alenezi et al. [17] introduced gradient and absorption differences associated with color channels into background light and transmission map estimation. Zhou et al. [23] introduced light attenuation features and feature prior in estimating background light and transmission map. Chandrasekar et al. [24] combine underwater image formation models and depth information to remove backscatter, and utilize lightweight networks and retinex model to further enhance the image. In addition, some methods introduced prior knowledge from other domains to improve visual quality. Berman et al. [25] used the haze-lines model to enhance the attenuated colors. Yuan et al. [7] enhanced the visibility of images by dark channel prior and color compensation in two color spaces, respectively. Due to the variability of underwater environments, the parameters and priors set by such methods are limited and cannot fully adapt to specific scenarios.
@@ 2.2 Deep Learning Methods
In recent years, deep learning methods have made significant progress in UIE tasks. Classical methods were devoted to optimize the network structure to achieve better visual enhancement. Jiang et al. [26] incorporated the laplacian pyramid to construct a lightweight multi-scale network, and improved the performance by recursive strategy and residual learning. Zhang et al. [27] introduced multi-level sub-networks to enrich the feature representation, and proposed a triple attention module to selectively cascade different features. To enrich the input features, some studies have introduced other color spaces. Wang et al. [28] introduced HSV space for the first time to adjust the brightness and saturation of underwater images. Zhou et al. [18] combined RGB and HSV color spaces to enhance the underwater images, and used attention and color-guided map to guide the network to perceive critical information. Furthermore, the researchers have delved into the structural composition of images from diverse perspectives of data distribution. By observing from the perspective of color spectrum distribution, Rao et al. [19] proposed a two-stage network using texture cues and color cues to enhance the underwater images. Wang et al. [29] integrated the wavelet transform and color compensation mechanism in the network to perform multi-level enhancement of details and colors.
To avoid the need for paired underwater images, unsupervised learning methods have received considerable attention. Among them, generative adversarial network (GAN) has been a prevalent framework. Jiang et al. [30] employed global-local adversarial mechanism to supervise image realism. Li et al. [31] introduced comparative learning into UIE and utilized the discriminator to evaluate the quality of different regions. Beyond GAN, other learning methods such as transfer learning and comparative learning also show immense potential. Jiang et al. [32] migrated image dehazing from air to the underwater domain, thus avoiding the reliance of paired underwater images. Liu et al. [11] constructed a twin adversarial contrastive learning framework and utilized underwater object detection tasks to guide image enhancement.
@@ 2.3 Depth-related Methods
The widespread application of depth information in image dehazing [33, 34, 35] has enlightened the understanding of underwater scenes in the UIE field. Traditional methods incorporated depth estimation to mitigate underwater degradation effects on image quality. Zhou et al. [36] used channel intensity prior and unsupervised method to generate two depth maps, further eliminated the blurring problem caused by scattering effects. Hou et al. [37] improved the estimation of transmission map by the local background light and reflectance map, which can eliminate the influence of non-uniform light in underwater scenes. Simultaneously, deep learning methods utilized depth information to guide the scene weights for targeted distribution. Chen et al. [20] introduced position attention and dilated convolution into the depth estimation network, guiding weight allocation by the predicted depth map. Fu et al. [12] reconstructed the raw image by predicting a depth-related transmission map, thereby obtaining the scene radiance for degradation. Cong et al. [13] used the depth map and the physical model for image restoration, with dual-discriminators constraining the generation of the restored image and the depth map.
@@ 3.1 Overall Architecture
As shown in Fig. 3, UVZ constitutes a two-stage UIE framework. In the first stage, we input the raw image $X\in\mathbb{R}^{H\times W\times 3}$ to the depth estimation network (DEN), and output the depth map $d\in\mathbb{R}^{H\times W\times 1}$ through encoding-decoding and two scale changes. Within the process of feature mapping, a dual-attention module (DAM) is introduced to monitor severely degraded regions from channel and spatial perspectives, allowing for a more explicit definition of the near and far regions. By introducing scene reconstruction and cyclic consistency loss, CycleGAN [38] ensures consistency between the target images and the source images. Inspired by this, we introduce the auxiliary supervision network (ASN) to generate the regression map $\widehat{X}\in\mathbb{R}^{H\times W\times 3}$ in the training phase, and jointly supervise the losses of DEN and ASN to promote greater correlation between $d$ and $X$. To refine the restoration of $\widehat{X}$, ASN fuses the decoded features from DEN via residual skip blocks (RSB), which further promotes the rationality of $\widehat{X}$ on the overall scene.
@@
In the second stage, we design a depth perception module (DPM) for near and far feature learning. DPM consists of two distinct branches for non-local and local feature learning, with selective fusion guided by $d$ from the previous stage. In response to the DPM, we design an R3S transformation for $d$, as shown in Fig. 4. Using the adjusted depth maps and the learnable parameters, the depth-guided enhancement network (DGEN) constructs a fusion strategy with receptive field adaptation to combine non-local and local features more rationally. Meanwhile, DGEN incorporates multiple residual blocks (RB) to solve the gradient vanishing. Both RB and RSB extract features through convolution-activation-convolution structures and residual connection, while the latter includes an additional 1 × 1 convolution block for fusion.
@@ 3.2 Depth Perception Module
We constructed a DPM that guides perception with depth, and adaptively enhance different regions through non-local and local learning branches. Specifically, to adapt the receptive fields of the two branches, the DPM applies an R3S transformation to the depth map $d$. As shown in Fig. 4, we first reshape d to $d$ to $d_{1}\in\mathbb{R}^{h\times w\times 1}$ consistent with the encoded feature scale, which is performed by nearest interpolation. Second, we alter the weight distribution of the far and near scenes by reversing the depth map $d_{rev}=1-d$, so that the non-local branch focuses more on high-quality information. In addition, we introduce a region smoothing (RS) process to remove noise from degraded regions, thereby avoiding the interference of far points with higher degradation on local information capture. RS divides the regions based on the size of the convolution kernel in the local branch, calculating the mean value of each region for normalization. The specific process is as follows:
@@
We adopt the Swin Transformer [39] to implement non-local branch for capturing dependencies. Compared to the Vision Transformer, Swin Transformer captures remote dependencies at diverse scales by constructing hierarchical feature mappings, which allows it to capture details of different shapes (e.g., fish, plants, and stones, etc.) in underwater scenes. Additionally, Swin Transformer introduces a shifted windows mechanism to enhance the receptive field, further extending the ability to facilitate non-local feature interactions. The input encoded features are first partitioned into a non-overlapping patch sequence $f_{i}=[p_{1},p_{2},\ldots,p_{m}]$, then flattened into linear vectors and computed as follows:
@@
$f_{ck}=d_{k}\times f_{ck}\times\beta_{k},k=1,3,5\$ (3)
@@ 3.3 Dual-Attention Module
In complex underwater imaging processes, the depth-related degradation attenuation is quite common. As shown in Fig. 5, we design a DAM that includes channel and spatial monitoring mechanisms to generate reasonable depth maps. First, DAM compresses the encoded feature $f$ into two channel descriptors through different pooling operations. Then, the channel dependence used to adapt the feature representation is captured by the squeeze-excitation process [40]. The above process can be represented as follows:
@@
where $AvgP(\cdot)$ and $AvgP(\cdot)$ represent the average pooling layer and maximum pooling layer, respectively, and $(\cdot)_{sig}$ is the sigmoid function. The features annotated with channel weights $f_{c}$ is input into the next stage, allowing regions with different degradation degrees to have distinct weights. First, we use the average pooling layer to reduce the feature scale, and then capture the spatial dependence by the query vector $q$, the key vector $k$, and the value vector $v$. In this process, we introduce a learnable parameter $\gamma$ to dynamically adjust the learning intensity:
@@
where $(\cdot)_{soft}$ is a softmax function, and $\gamma$ is a learnable parameter used to adjust the intensity. We use the deconvolution $(\cdot)_{dc}$ to reshape features into input scale and annotate spatial weights $f_{s}$ back to input features. We fused the two attention features and perform residual connections with the original input $f$. As a result, the proposed model can capture the hidden feature $f_{h}$ with regional degradation prominence:
$f_{h}=f+f_{c}\times\left(f_{s}\right)_{dc}$ (7)

### v4

@@ 1 Introduction
As a crucial information carrier, underwater images play an important role in ocean tasks(Zhang et al., 2024b). Unfortunately, underwater images often encounter various degradation problems due to two main factors (Park and Eom, 2024). Firstly, different wavelengths of light decay at varying rates in the underwater medium, resulting in a prevalent blue-green color deviation. Secondly, the scattering effect of suspended particles hinders normal light propagation, leading to problems such as low contrast and blurred details (Zhang et al., 2022). Meanwhile, these degradations worsen with the increasing distance from the imaging scene to the camera (Guo et al., 2019). Therefore, underwater image enhancement (UIE) methods are widely used in visual interpretation disciplines, such as biology, ecology, and animal behavior, aiding researchers in conducting subsequent analysis. Meanwhile, as a pre-processing step for computer vision tasks such as image segmentation, saliency detection, and moving object detection, UIE methods can significantly improve the accuracy and robustness of algorithms, providing strong support for ocean monitoring, ecological protection, and resource exploration.
Figure 1: Near-far scenarios enhancement comparison by different methods, where the orange box represents the enlarged near scene and the red box represents the enlarged far scene. The top row shows the raw image, the results of L2UWE (Marques and Albu, 2020), CBM (Yuan et al., 2020), TEBCF (Yuan et al., 2022), WWPE (Zhang et al., 2024c), and Water-Net (Li et al., 2019). The bottom row shows the results of SCNet (Fu et al., 2022b), TACL (Liu et al., 2022), USUIR (Fu et al., 2022a), PUGAN (Cong et al., 2023), Spectroformer (Khan et al., 2024), and our UVZ.
In the early days of UIE field, researchers have proposed model-free and model-based traditional methods. Model-free methods (Marques and Albu, 2020; Ancuti et al., 2018; An et al., 2024) enhances underwater images by modifying pixels in degraded channels and blurred regions. On the other hand, model-based methods (Yuan et al., 2022; Alenezi et al., 2022) introduce physical cues to build the underwater degradation process, and reverse it to restore underwater images. In recent years, deep learning methods (Liu et al., 2022; Zhou et al., 2023a; Rao et al., 2024) have shown remarkable enhancements, leveraging benchmark data and miscellaneous networks. Despite significant progress achieved with the above methods, challenges emerge when addressing widely varying degrees of regional degradation. As shown in Figure 1, most UIE methods have considerable visual improvement in the near scene, but still retain notable color deviation in the far scene. Figure 2 highlights the significant differences between near and far scenes, thereby underscoring the motivation behind this work. A more detailed description is provided in Section 3.1.
@@
For more precise color enhancement, we believe that the integration of depth information related to scene distance is essential, which benefits the region division in underwater images. Previous works (Fu et al., 2022a; Cong et al., 2023; Chen et al., 2023) attempted to incorporate depth information related to scene distance, contributing to the response of degraded regions. However, direct weight adjustments may restrict the exploration of visual cues, as the adjacent features of far and near points exhibit varying gains in the enhancement process. In near regions of relatively good quality, local information from surrounding features is critical to the quality of the enhanced image, whereas in the far regions of heavily degradation, non-local information from other regions can provide more valuable insights. With the above analysis, we aim to construct an adaptive region enhancement method in underwater scenes, achieving comprehensive visual enhancement by differentiating in utilizing local and non-local information. To achieve this goal, there are two main challenges:(1) On the one hand, the local feature capture capability helps to enhance the near region where the surrounding information is enough. On the other hand, the remote dependency establishment capability helps to enhance the far region where the surrounding information is lost. However, due to feature space differences, there is mutual interference between these two capabilities. (2) Diverse underwater scenes tremendously raise the difficulty of predicting the depth map.
@@
• We design a novel depth-guided perception framework, UVZ, for non-homogeneous degraded underwater scenes, which utilizes adapted depth maps to execute a learning strategy that integrates non-local and local features.
• To improve the distinction of degraded regions, we incorporate an attention estimation-scene reconstruction training process, aiming to strengthen the connection between depth maps and underwater scenes.
• Compared to state-of-the-art methods, UVZ exhibits more attractive results across extensive qualitative and quantitative comparisons, and demonstrates generalization abilities to other visual tasks without parameter tuning.
@@ 2.1 Traditional Methods
In response to the degradation characteristics of images, model-free methods notably enhance the contrast and saturation of the images. The early studies focused on adjusting image pixel values in the spatial domain. Ancuti et al. (Ancuti et al., 2018) fused color compensated and white balance images to obtain natural underwater image. Marques et al. (Marques and Albu, 2020) used contrast to guide the generation of images with salient details and brightness enhancement. As the research progressed, some methods introduced morphology and structural properties to understand underwater images. Based on the concept of contour bougie morphology, Yuan et al. (Yuan et al., 2020) enhanced the image quality by mitigating scattering blur and contrast stretching. Kang et al. (Kang et al., 2023) decompose the image into mean intensity, contrast, and structure, the three components are fused in different ways to obtain the enhancement result. Additionally, Zhang et al. (Zhang et al., 2022) introduced the attenuation characteristics to correct the color, and introduced the integral and squared integral maps to enhance the contrast. Overall, model-free methods overlook the characteristics of underwater degradation process, especially the depth between the object and the camera (Zhang et al., 2023; Zhou et al., 2023c), leading to color distortion or improper enhancement.
Model-based methods explore the process of underwater image degradation and use priors and assumptions to model this process. Alenezi et al. (Alenezi et al., 2022) introduced gradient and absorption differences associated with color channels into background light and transmission map estimation. Zhou et al. (Zhou et al., 2023c) introduced light attenuation features and feature prior in estimating background light and transmission map. Chandrasekar et al. (Chandrasekar et al., 2024) combine underwater image formation models and depth information to remove backscatter, and utilize lightweight networks and retinex model to further enhance the image. In addition, some methods introduced prior knowledge from other domains to improve visual quality. Berman et al. (Berman et al., 2021) used the haze-lines model to enhance the attenuated colors. Yuan et al. (Yuan et al., 2022) enhanced the visibility of images by dark channel prior and color compensation in two color spaces, respectively. Due to the variability of underwater environments, the parameters and priors set by such methods are limited and cannot fully adapt to specific scenarios.
@@ 2.2 Deep Learning Methods
In recent years, deep learning methods have made significant progress in UIE tasks. Classical methods were devoted to optimize the network structure to achieve better visual enhancement. Jiang et al. (Jiang et al., 2022a) incorporated the laplacian pyramid to construct a lightweight multi-scale network, and improved the performance by recursive strategy and residual learning. Zhang et al. (Zhang et al., 2024a) introduced multi-level sub-networks to enrich the feature representation, and proposed a triple attention module to selectively cascade different features. To enrich the input features, some studies have introduced other color spaces. Wang et al. (Wang et al., 2021) introduced HSV space for the first time to adjust the brightness and saturation of underwater images. Zhou et al. (Zhou et al., 2023a) combined RGB and HSV color spaces to enhance the underwater images, and used attention and color-guided map to guide the network to perceive critical information. Furthermore, the researchers have delved into the structural composition of images from diverse perspectives of data distribution. By observing from the perspective of color spectrum distribution, Rao et al. (Rao et al., 2024) proposed a two-stage network using texture cues and color cues to enhance the underwater images. Wang et al. (Wang et al., 2024) integrated the wavelet transform and color compensation mechanism in the network to perform multi-level enhancement of details and colors.
To avoid the need for paired underwater images, unsupervised learning methods have received considerable attention. Among them, generative adversarial network (GAN) has been a prevalent framework. Jiang et al. (Jiang et al., 2022c) employed global-local adversarial mechanism to supervise image realism. Li et al. (Li et al., 2023) introduced comparative learning into UIE and utilized the discriminator to evaluate the quality of different regions. Beyond GAN, other learning methods such as transfer learning and comparative learning also show immense potential. Jiang et al. (Jiang et al., 2022b) migrated image dehazing from air to the underwater domain, thus avoiding the reliance of paired underwater images. Liu et al. (Liu et al., 2022) constructed a twin adversarial contrastive learning framework and utilized underwater object detection tasks to guide image enhancement.
@@ 2.3 Depth-related Methods
The widespread application of depth information in image dehazing (Liu et al., 2018; Ju et al., 2021; Yang et al., 2022) has enlightened the understanding of underwater scenes in the UIE field. Traditional methods incorporated depth estimation to mitigate underwater degradation effects on image quality. Zhou et al. (Zhou et al., 2023b) used channel intensity prior and unsupervised method to generate two depth maps, further eliminated the blurring problem caused by scattering effects. Hou et al. (Hou et al., 2024) improved the estimation of transmission map by the local background light and reflectance map, which can eliminate the influence of non-uniform light in underwater scenes. Simultaneously, deep learning methods utilized depth information to guide the scene weights for targeted distribution. Chen et al. (Chen et al., 2023) introduced position attention and dilated convolution into the depth estimation network, guiding weight allocation by the predicted depth map. Fu et al. (Fu et al., 2022a) reconstructed the raw image by predicting a depth-related transmission map, thereby obtaining the scene radiance for degradation. Cong et al. (Cong et al., 2023) used the depth map and the physical model for image restoration, with dual-discriminators constraining the generation of the restored image and the depth map.
@@ 3.1 Overall Architecture
As shown in Figure 3, UVZ constitutes a two-stage UIE framework. In the first stage, we input the raw image $X\in\mathbb{R}^{H\times W\times 3}$ to the depth estimation network (DEN), and output the depth map $d\in\mathbb{R}^{H\times W\times 1}$ through encoding-decoding and two scale changes. Within the process of feature mapping, a dual-attention module (DAM) is introduced to monitor severely degraded regions from channel and spatial perspectives, allowing for a more explicit definition of the near and far regions. By introducing scene reconstruction and cyclic consistency loss, CycleGAN (Zhu et al., 2017) ensures consistency between the target images and the source images. Inspired by this, we introduce the auxiliary supervision network (ASN) to generate the regression map $\widehat{X}\in\mathbb{R}^{H\times W\times 3}$ in the training phase, and jointly supervise the losses of DEN and ASN to promote greater correlation between $d$ and $X$. To refine the restoration of $\widehat{X}$, ASN fuses the decoded features from DEN via residual skip blocks (RSB), which further promotes the rationality of $\widehat{X}$ on the overall scene.
@@
In the second stage, we design a depth perception module (DPM) for near and far feature learning. DPM consists of two distinct branches for non-local and local feature learning, with selective fusion guided by $d$ from the previous stage. In response to the DPM, we design an R3S transformation for $d$, as shown in Figure 4. Using the adjusted depth maps and the learnable parameters, the depth-guided enhancement network (DGEN) constructs a fusion strategy with receptive field adaptation to combine non-local and local features more rationally. Meanwhile, DGEN incorporates multiple residual blocks (RB) to solve the gradient vanishing. Both RB and RSB extract features through convolution-activation-convolution structures and residual connection, while the latter includes an additional 1 × 1 convolution block for fusion.
@@ 3.2 Depth Perception Module
We constructed a DPM that guides perception with depth, and adaptively enhance different regions through non-local and local learning branches. Specifically, to adapt the receptive fields of the two branches, the DPM applies an R3S transformation to the depth map $d$. As shown in Figure 4, we first reshape d to $d$ to $d_{1}\in\mathbb{R}^{h\times w\times 1}$ consistent with the encoded feature scale, which is performed by nearest interpolation. Second, we alter the weight distribution of the far and near scenes by reversing the depth map $d_{rev}=1-d$, so that the non-local branch focuses more on high-quality information. In addition, we introduce a region smoothing (RS) process to remove noise from degraded regions, thereby avoiding the interference of far points with higher degradation on local information capture. RS divides the regions based on the size of the convolution kernel in the local branch, calculating the mean value of each region for normalization. The specific process is as follows:
@@
We adopt the Swin Transformer (Liu et al., 2021) to implement non-local branch for capturing dependencies. Compared to the Vision Transformer, Swin Transformer captures remote dependencies at diverse scales by constructing hierarchical feature mappings, which allows it to capture details of different shapes (e.g., fish, plants, and stones, etc.) in underwater scenes. Additionally, Swin Transformer introduces a shifted windows mechanism to enhance the receptive field, further extending the ability to facilitate non-local feature interactions. The input encoded features are first partitioned into a non-overlapping patch sequence $f_{i}=[p_{1},p_{2},\ldots,p_{m}]$, then flattened into linear vectors and computed as follows:
@@
$f_{ck}=d_{k}\times f_{ck}\times\beta_{k},k=1,3,5.\$ (3)
@@ 3.3 Dual-Attention Module
In complex underwater imaging processes, the depth-related degradation attenuation is quite common. As shown in Figure 5, we design a DAM that includes channel and spatial monitoring mechanisms to generate reasonable depth maps. First, DAM compresses the encoded feature $f$ into two channel descriptors through different pooling operations. Then, the channel dependence used to adapt the feature representation is captured by the squeeze-excitation process (Hu et al., 2018). The above process can be represented as follows:
@@
where $AvgP(\cdot)$ and $MaxP(\cdot)$ represent the average pooling layer and maximum pooling layer, respectively, and $(\cdot)_{sig}$ is the sigmoid function. The features annotated with channel weights $f_{c}$ is input into the next stage, allowing regions with different degradation degrees to have distinct weights. First, we use the average pooling layer to reduce the feature scale, and then capture the spatial dependence by the query vector $q$, the key vector $k$, and the value vector $v$. In this process, we introduce a learnable parameter $\gamma$ to dynamically adjust the learning intensity:
@@
where $(\cdot)_{soft}$ is a softmax function, and $\gamma$ is a learnable parameter used to adjust the intensity. To label spatial weights $f_{s}$ on input features, we first use a deconvolution $(\cdot)_{dc}$ to reshape $f_{s}$ into input scale. Subsequently, we fuse the two attention features and perform residual connections with the original input $f$. As a result, the proposed model can capture the hidden feature $f_{h}$ with regional degradation prominence:
$f_{h}=f+f_{c}\times\left(f_{s}\right)_{dc}.$ (7)